Sodankylä
Geophysical
Observatory
Oulun yliopisto

NEMI Campaign 2008
Diary

Last update: 2015-08-28, 11:28:55 UTC.


2008-02-08

09:30-11:30 Heating/VHF. Quiet.

Kilpisjärvi: The forecast made by YR seems to be good. They promised clear sky from 20UT and now the sky opens. Also we start to see weekend effect in KIL ASC is visible. Norwegians arrived and snow mobile rally on the ice of the lake Kilpisjärvi has started already. Already lights of village are blocked from the E-S sector.


2008-02-07

Heating and VHF experiment in daytime: No D-layer today. Possibly a very weak sporadic E in the beginning.

Evening: Heating stopped early due to low activity in the evening. The radar changed to FPI mode over KEOPS already before 20 UT. Tromsø cleared up after rain and snow, and there was bright aurora from around 23:50. RSC running from about 19 UT, and ODIN as well.

Kilpisjärvi: After couple of cloudy nights we had clear sky from the sun set until 23UT. 2325UT auroras activated but the sky was not clear anymore. MSP is now pointed to KEOPS volume and also spectrometer turned to new direction.


2008-02-06: The VHF experiment started late today due to a failing cooling pump. It seems that water pumps do not like temperature changes. Also this VHF run was combined with Heating; first Heating was done in 6 sec on, 6 sec off, X-mode. Due to a strong sporadic E layer, this was changed to 3 min on, 3 min off, X-mode at 1012 UTC.

Alex has taken the ODIN camera to the Heating workshop. As it turns out he is a professional TV repairman who has turned to studying electronic engineering. ODIN is now mounted indoors in one of the PRE domes. It gives output but for some reason the sensitivity is very low.

UHF runs in the night: Clear sky and auroral activity. The HotArc heating experiment was run 18-22:30 UT, combining 5 Hz and CW modulation. The FPI stations were cloudy so the UHF tristatic run continued over Tromsø until end of scheduled time. RSC and SBIG imaging in magnetic zenith.


2008-02-05: Another clear morning in Tromsø. The water pump fails again.

The ionosphere is weaker today but there is a PMWE layer. We have exchanged some VHF hours for heating and are running a VHF-Heating PMWE modulation experiment (data).

Moving data from the SBIG computer to the external disk. The internal disk is 40 GB, which is really small when taking images with intervals of a few seconds.

Kilpisjärvi (1735UT): The sky is cloudy and the forecast does not look good. Huge signal from scattered lights of village. Esa continued trip towards Sodankylä. Today we told about HotPay2 rocket at Kilpisjärvi school and we took part to school's Laskiaisrieha.

Tromsø: Carl-Fredrik picked up Björn from the bus stop at Fagernes. Heating experiment (hot_arc) 18-22 UTC; very quiet conditions however. UHF measuring tristatic tau2pl field-aligned over Tromsø all night (18-01 UTC). No data from the FPIs due to clouds above Kiruna.


2008-02-04: Antti and Timo ran the VHF experiment during daytime (data). Carl-Fredrik, Esa, Anasuya and Eoghan went to the airport via Blå Rock. Anasuya and Eoghan picked up a rental car and drove Esa to Skibotn where Tero waited to take him to Kilpis. Alexander Witte, another German student who left for Scandinavia and is now making a diploma project building an ionosonde antenna for KTH in central Stockholm, arrived at the same time and was taken to Eiscat by Carl-Fredrik. He has already helped sorting out some technical problems here.

Alexander and Björn installed the ODIN camera at PRE. The night was clear in TRO with some weak stable arcs. Kiruna was clear in the evening but cloudiness increased during the night. The other stations were at least partly cloudy. UHF tristatic run at 240 km all night.

The scanning photometer was turned towards Kiruna (azimuth 158 degrees), at least roughly. A scan for Bellatrix, Betelgeuze and Mars has been programmet in order to check the exact direction. However, it now stands on top of ice ridges. Hope these will melt until next night.

Kilpisjärvi: The MSP turned towards KEOPS volume but no clear sky to confirm the direction. The sky stayed cloudy whole evening. Later on night the sky came clear and some pulsation auroras were on the sky between 02-03UT. Esa spent night in Kilpisjärvi first time during this century and got taste of site station life. I also showed IRIS to him when we arrived from Skibotn to Kilpisjärvi.


2008-02-03: VHF run with some PMWE during the day, run by Antti (data).

Anasuya and Eoghan from UCL arrived for the tristatic Eiscat drift velocity vs FPI neutral wind velocity comparison above Kiruna. UHF tristatic measurements started at 240 km for the bistatic red-line FPIs at Sodankylä and Esrange, switching to 110 km for the Esrange green-line FPI.

Esa returned with a late evening flight and showed videos from Andøya.


2008-02-02: The midday VHF experiment of today is two hours longer due to scheduling requirements on the weekend. This turns out to be fantastic, because we have a "majestetic D layer" (Antti) as well as plenty of PMWE echos, sometimes two simultaneous layers. What a pity we cannot quickly turn on the Heater! This is what the data looks like.

Tero reports cloudy sky and snowfall from KIL this morning.


2008-02-01: All quiet after the launch. The experiment was largely successful: at least the telemetry worked and the data is stored. We have to wait for the analysis. More precise information also on technical matters like the exact trajectory will come on Friday, 9th February if all goes well.

Water at the EISCAT site in Tromsø seems still to be unstable. The VHF midday experiment ran without problems. During the run a few phone calls were made between our team and Cesar La Hoz, who is the PI of the MORRO radar. He informed us, that he runs MORRO whenever there's no VHF in the schedule. Our VHF try last night was not scheduled. He also said that one can call him any time and ask for MORRO to be stopped, but we did not know that. Ok, next time...

In the evening, after the CP run ended, we still had two hours on the UHF. Thanks to our British colleagues, from now on we run tristatic. We use the tau2pl experiment with a common volume 240 km above KEOPS in support of the UCL FPIs. However, the weather was cloudy at Tromsø and above Kiruna.

Esa enjoys the launch party at Andøya in the evening. Kippis!


2008-01-31: From Esa: Thursday morning opened blue skies partly both in Tromso and at Andoya. Weather forecast today is best so far and the aurora might appear tomorrow if not already today. Last night however was the quietest one could expect! A perfect reference night for NEMI. EISCAT VHF worked well today.

In the evening the UHF run was taken care of by EISCAT within the CP pool. They ran tristatic beata. We tried to start the VHF when it became clear that the rocket will fly tonight, but this was impossible due to strong interference from the MORRO MST radar.

Rocket was launched just before 19:14 UTC! A little movie clip of the mood during the launch night is available here. Also, there are photos of the NEMI aurora seen from Kilpisjärvi as well as the official launch photos from Andøya Rocket Range.

From Tromsø: There was indeed some blue sky in the morning. Antti running the VHF arc_dlayer_ht daytime experiment. After some problems with the timing and disk space at EISCAT we got 15 extra minutes after the scheduled end of run. The UHF is running CP beata. The night-time UHF measurement will also be taken care of by the CP.

Kilpisjärvi: We had clear sky most of the evening. T-1h we had active auroras in the zenith. Arc started to travel north and RSC (Aurora TV) had the arc in the FOV during the launch. Afterwards, 20UT, we got nice display northern half of the sky was filled by auroras. At 21UT we started to get some frost fog above the village and lake. See photos.


2008-01-30: From Esa: Wednesday morning showed blue skies over Andøya Rocket Range at 08:00 in the morning! However heavy clouds were drifting from the sea towards the continent. But the winds have calmed down below the limit of 12 m/s, which is necessary for a possible launch. The weather forecast however now maybe favours Friday as a most probable clear night.

I think we are expecting perfect launch conditions for NEMI, due to the very weak auroral activity during the recent 3 nights, especially last night. If the coronal hole induced activity hits us on the coming days, plus the weather gets better, plus the launch of HotPay II is successful within the planned criteria, we would have a measurement of a fairly undisturbed air during upleg and ending at downleg with actual measurement of air hit by precipitating electrons.

From Tromsø: Some decimetres of wet snow in the morning, which was good, because the pump was broken again and we melted snow for coffee. CF dug the optical instruments out from piles of snow. There are in fact also some patches of blue sky. Daytime runs: VHF arc_dlayer_ht and CP UHF beata. Reimei passage in the night, which would be ideal in conjunction with the rocket. The sky may be partly clear tonight so at least we could turn the MSP. Update: more snow forecasted for tonight.

Timo ran the UHF in the night, monostatic arc1 from 19UT changing to SW time and adding the remotes at 23UT for the ALIS-Reimei conjunction. The VHF was running northward observations for Cluster (AA programme). Heavy snowfall all night. Björn checked the ODIN camera and brought it back to the university since it did not work. Tima Sergienko ran ALIS from Kiruna. 19-23 we tested the new Andoya position and a filter sequence synchronised with the Kilpisjärvi and Sodankylä cameras.

From Esa: Last night's countdown was cancelled at 3 hrs before the end of the nominal launch window. The weather radar and local forecasts showed that the snowstorm conditions would prevail for more than 1 hour after the decision was made, and at that time none of the launch criteria were fulfilled. The countdown was on hold at -02:00 hrs and it should have been restarted latest at 2 hrs before the end of the window. The actual shortest possible time to launch after the hold situation would be 1hr 30min - 1hr 45min, if a fast launch would be needed. So yesterday both the operation manager and PI saw that a launch will not be possible with the weather conditions. For us this was a lucky situation, because we also want some aurora!


2008-01-29:

Launcher photo by Andøya Rocket Range.

From Esa: HotPay II countdown will start on Tuesday 29th of January at 1600 UTC, 3 hrs before the nominal launch window at 1900 UTC. The countdown is a complicated procedure, starting with the first weather balloon launch, informing the ATC, alerting the fire fighting service and ambulance etc. First possible hold is at -2 hrs for weather, next one at -30 minutes for science conditions. Next hold (for science) may occur at 12 mins prior to the launch and a short term hold for science at -3 mins. During the last 60 seconds final GO decisions by the PI and 5 operational units are checked so that a final GO/ NO GO status is announced. The final countdown is of course the last 10 seconds to T-0. Countdown is continued at T+time until ballistic impact.

Last nights auroral activity was much weaker than the weak activity during the night before. It does not yet rain whales at Andoya, but the wind is strong, 20 m/s, and we get continuous rain, hails and wet snow. The wind probably does not allow a launch today. More news later.

Kilpisjärvi: Tero arrived to Kilpisjärvi. We have meridian scanning photometer (later MSP), real speed camera (RSC), AP7-digital camera and spectrometer in operation. Also FMI's new EMCCD all-sky imager is in operation. We had some aurora activity in Kilpisjärvi. Arc rose to zenith and some active aurora in NW direction before 24UT.

Tromsø: Snowing all day. VHF arc_dlayer_ht experiment in the day and UHF beata in the night. Eiscat was also running UHF beata in the daytime. The optical instruments were covered with snow all the time, so we got no useful observations here.


2008-01-28: Beautiful cloudy day here in TRO, a few degrees below 0C, 30 cm fresh snow on the yard - and no water in Hilton nor the main building! We have to melt snow for drinking water.

Update: It seems the pump is fixed - water again! The system is in the PRE building and CFE took a look at it. Working showers and WCs are a luxury we need.

We drove Esa to the airport and he arrived safely at Andenes. Tero left in the afternoon and drove the Passat to Kilpis.

CFE bought cables for connecting Tero's old laptop to the VHS soundtrack. Björn has installed ODIN, but the weather did not allow tests. Since the optical campaign has started, the weather has changed into heavy rain. (After all, this is Tromsø). We plan to turn the three MSPs at KAR, KIL and TRO to scan in the same plane over Eiscat, but would need visible bright stars for that.

Late in the evening at the very end of this day Timo P turned up with the Oulu VW bus, after dropping off instruments at Kilpisjärvi.


2008-01-27: Esa and Antti (Passat) as well as Carl-Fredrik and Tero (Land Cruiser) drove from Sodankylä to Tromsø, experiencing extremely variable weather: starting at Sodankylä from below -20C, up to convenient 0C at Tromsø, meeting however heavy snowfalls occasionally on the road. The start was delayed by one hour due to the need of fixing the driver's door of the Land Cruiser, after a minor unexpected collision. Luckily we got our technician Pertti on Sunday to fix the door. In addition to the planned stops at sites in Finland, Tero and Carl-Fredrik encountered a short-lasting unexpected road close in Skibotn, due to rescue of a 4-wheeler with trailer, which had slipped off the road.

The friendly face of the Tromsø city - the white pearl in the north - rewarded the travellers who found excellent hamburgers and a short moment of rest in a local pub called Blå Rock. As usual, the faces of the Tromsø people looked friendly, too.

At EISCAT Hilton we faced the normal scientist work: clean up the parking places for the cars and make a path to Hilton's front door - in the 20 cm of fresh snow! Now at 11:30 pm local time unloading and installing the optical instrumentation is still being done by Tero and Carl-Fredrik.

A phone call to Jacek by Carl-Fredrik: NEMI is installed on the payload at Andenes, after some tedious work by Misha and Jonas. Tomorrow Jacek has the same flight to Andenes as Esa.

Some very weak auroral activity in the north between 1930-0100 UTC, with momentary brightenings of the oval.


2008-01-26: A day of rest before the storm. No campaign activities today.


2008-01-25: Timo collects keys for VW van from U Oulu.


Last modified: 28 August 2015, 11:28:55.